### v2

# Cold-Start Recommendation with Knowledge-Guided Retrieval-Augmented Generation

Abstract.
Cold-start items are a long-standing challenge for recommender systems because they lack the historical user interactions required by collaborative models. Recent zero-shot approaches employ large language models (LLMs) to avoid this dependence, yet they still struggle with sparse metadata and limited or hallucinated knowledge. We introduce ColdRAG, a retrieval-augmented generation framework that equips an LLM with a dynamically built, domain-specific knowledge graph so it can deliver accurate recommendations in cold-start settings without task-specific fine-tuning. ColdRAG first transforms any available structured item fields into a rich natural-language profile. It then extracts entities and relations from each profile and integrates them into a shared knowledge graph that captures item semantics through explicit connections. Finally, given a user’s interactions as a query, it traverses the graph using LLM-guided edge scoring, retrieves candidate items along with supporting evidence, and prompts the LLM to rank them for recommendation. Through multi-hop reasoning over the constructed knowledge graph, ColdRAG overcomes the sparse-metadata bottleneck by grounding its recommendations in verifiable evidence, substantially reducing hallucinations and exploiting semantic connections among items. Experiments on three public benchmarks show that ColdRAG outperforms state-of-the-art zero-shot baselines in both Recall and NDCG. By unifying knowledge-graph reasoning with retrieval-augmented LLM generation, ColdRAG provides a practical solution to the real-world cold-start recommendation problem.
Recommendation System, Retrieval-Augmented Generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Recommender systems underpin e-commerce, streaming, and advertising (Sarwar et al., 2001; Koren et al., 2009), yet performance in the cold-start setting—where new users or items have little to no interaction history—remains fragile (Gope and Jain, 2017; Zhang et al., 2025). Collaborative filtering (CF) depends on dense user–item signals and degrades sharply under sparsity. Content-based (CB) approaches leverage side information (e.g., item attributes, user profiles) but are constrained by incomplete or noisy metadata (Lam et al., 2008; Volkovs et al., 2017b; Gu et al., 2016). Hybrid models (CF+CB) partly alleviate these issues, yet still require some interactions to anchor predictions (Volkovs et al., 2017a; Zhang and Wang, 2015). Consequently, robust cold-start recommendation remains a central obstacle as catalogs and user populations evolve continuously.
Recent advances in large language models (LLMs) have opened a zero-shot avenue to recommendation by casting it as natural-language generation and prompting the model to directly infer relevant items (Kojima et al., 2022; Hou et al., 2024; Li et al., 2019; Xu et al., 2025). Because the cold-start setting is inherently zero-shot, a growing body of work explores LLM-based strategies for new users and items, leveraging pretrained world knowledge with little or no domain-specific data (Li et al., 2019; Hou et al., 2024). In particular, prompting LLMs to score or rank candidates from textual cues—optionally aided by lightweight keyword extraction or retrieval—achieves competitive cold-start performance without fine-tuning (Hou et al., 2024; Liang et al., 2025; Kieu et al., 2024). Complementary approaches organize candidates hierarchically and iteratively prune them with tree-style reasoning to sharpen recommendations (Zhang et al., 2024). Despite this promise, prompt-only LLM recommenders remain vulnerable to hallucination, stale or niche inventory coverage gaps, and generation instability, underscoring the need for retrieval augmentation and robustness techniques (Huang et al., 2025b; Jiang et al., 2025; Islam and Faruk, 2025; Huang et al., 2025a; Holtzman et al., 2020; Yang et al., 2024).
Table 1. Comparison of recommender system categories. CF: Collaborative Filtering (based on interaction data); CB: Content-Based (uses item/user attributes). Each column represents a key capability: support for cold-start recommendation, grounding in external evidence, zero-shot usability, and multi-hop reasoning over knowledge graphs. (* denotes static metadata, not retrieved evidence.)
Cold-start Grounding Zero-shot Multi-hop
CF
CB *
CF + CB
Prompt-only LLM
RAG-based LLM
ColdRAG (ours)
Retrieval-augmented generation (RAG) grounds outputs in external evidence (Lewis et al., 2020; Peng et al., 2024), but common variants misalign with cold-start setting. Similarity-embedding RAG flattens structure and returns top-$k$ textually similar items that may be semantically adjacent yet decision-irrelevant. Naive GraphRAG pipelines inject structure but either rely on global community summaries—useful for topic-level queries yet too coarse for item-level personalization—or perform $k$-hop expansion that sweeps in all neighbors within $k$ hops (Xu et al., 2024; Edge et al., 2024; Guo et al., 2024), inducing exponential neighborhood growth that quickly exhausts context and dilutes the few relations that justify a cold-start choice (Xu et al., 2018; Topping et al., 2021; Lee et al., 2025). In cold-start, what matters is not neighborhood size but which relations and paths (e.g., shared genres, themes, creators) carry explanatory power (Kieu et al., 2024).
We target an LLM-augmented, KG-based recommender that is robust to item cold-start settings. To this end, we pursue three design goals: Knowledge grounding & generation stability—constrain outputs to KG-retrieved candidates with explicit evidence, reducing hallucinations and stabilizing outputs; Multi-hop Reasoning-based Retrieval—prioritize semantically meaningful relation paths over bulk neighborhoods to form a compact, high-quality candidate set; and Cold-start adaptability with explainability—integrate new items into the KG for immediate recommendations, accompanied by concise, path-based rationales.
In this work, we propose ColdRAG that casts retrieval as goal-directed path selection: during traversal, relations are scored for how much they advance the recommendation objective, and branches that add little explanatory value are pruned early. This converts KG structure from a passive index into an active reasoning space that builds compact evidence chains under a fixed token budget. ColdRAG proceeds in four stages: (i) Item profile generation converts available fields (titles, attributes, reviews) into coherent natural-language profiles; (ii) Knowledge graph construction extracts typed entities and relations to form a domain-specific KG; (iii) Multi-hop reasoning-based retrieval scores relation paths and prunes unrelated branches to retrieve a compact candidate set with supporting evidence; and (iv) Recommendation with retrieval-augmented generation ranks candidates using associated supporting facts to produce a concise, source-cited list. By grounding generation in explicit evidence and constraining traversal to high-yield paths, ColdRAG mitigates hallucination, incorporates new items seamlessly, and improves stability—without incurring exponential neighborhood expansion.
We highlight the advantages of ColdRAG in Table 1, which shows representative categories (CF/CB/Hybrid, prompt-only LLMs, and RAG variants) along four axes: Cold-start, Grounding, Zero-shot, and Multi-hop. In summary, our contributions are threefold:
  • We diagnose why prompt-only LLM recommenders and conventional RAG (similarity- and $k$-hop–based) underperform in cold-start, motivating structured reasoning over explicit knowledge.
  • We propose ColdRAG, integrating retrieval-augmented generation with dynamically constructed KG reasoning and path-selective, budget-aware retrieval that returns evidence chains without task-specific training.
  • We demonstrate state-of-the-art zero-shot results on multiple benchmarks and analyze the stability gains attributable to path-selective retrieval and evidence-grounded generation.
2. Related Works
Figure 1. Overview of the proposed ColdRAG framework. Given raw metadata for each item, a LLM generates a natural language item profile that denoises and enriches key concepts. From these profiles, structured entities and relations are extracted and used to format a semantic knowledge graph. During inference, ColdRAG retrieves relevant candidate items and contextual knowledge by performing LLM-guided multi-hop reasoning over the graph based on a user’s interaction history. These are then assembled into prompts and passed to the LLM, which generates the final recommendation list in a zero-shot manner.
2.1. Cold-Start Recommendation
The cold-start problem (Zhang et al., 2025) is a well-recognized challenge in recommendation systems, occurring when new users or items lack sufficient historical interaction data, prompting early content-based or hybrid techniques that leverage auxiliary information—such as item descriptions or user attributes—to compensate for missing collaborative signals. Representative approaches include collaborative deep learning, which jointly embeds textual content and interaction data (Wang et al., 2015), and meta-learning frameworks that treat cold-start as a few-shot learning problem (Lee et al., 2019), reflecting a broader trend toward transferring knowledge from side information or related tasks. More recent efforts add explicit cold-start mechanisms: DropoutNet simulates cold-start by randomly dropping out warm-item embeddings during training to learn robust representations (Volkovs et al., 2017a), while contrastive learning methods like CLCRec preserve collaborative signals in content-derived embeddings to improve new-item recommendations (Wei et al., 2021). Distributional robustness offers another promising direction, with TDRO applying distributionally robust optimization to account for temporal shifts in item features and thereby enhance generalization (Lin et al., 2024). General sequential recommenders like SASRec (Kang and McAuley, 2018) that rely on item IDs tend to struggle on unseen items, a limitation addressed by UniSRec (Hou et al., 2022) which replaces IDs with textual embeddings to enable cross-domain transfer. Still, these models either require dedicated cold-start modules, extensive side information, or retraining when new items emerge, whereas our proposed ColdRAG framework avoids finetuning by leveraging an LLM with retrieval augmentation to enable zero-shot recommendations through generative reasoning.
2.2. LLM-based Recommender Systems
Early efforts such as P5 treat diverse recommendation tasks as a unified text-to-text problem, fine-tuning a T5 backbone to produce recommendations, explanations, and other outputs with a single model (Geng et al., 2022). Subsequent studies exploit the semantic richness of pre-trained language encoders to embed item descriptions, reviews, or user profiles, improving data efficiency and cross-domain transfer. With the advent of powerful LLMs, researchers have explored prompt-based, zero-shot recommenders. LLMRank casts ranking as instruction following: given a user’s history and candidate descriptions, an LLM ranks items without additional training (Hou et al., 2024). PromptRec shows that carefully designed prompts enable even relatively small language models to handle extreme cold-start settings (Wu et al., 2024). TaxRec augments prompts with taxonomy tokens so the LLM can better navigate hierarchical item spaces (Liang et al., 2025). KALM4Rec couples a graph-based keyword retrieval module with prompt-based LLM re-ranking for cold-start user recommendation without fine-tuning (Kieu et al., 2024). While these models demonstrate impressive zero-shot ability, they rely solely on the prompt’s immediate text and thus can hallucinate or miss domain-specific details. Our ColdRAG differs by explicitly retrieving and supplying relevant item knowledge—structured as a dynamically built graph—to the LLM, grounding its reasoning and improving accuracy for newly introduced items.
3. Proposed Method
We propose a Retrieval-Augmented Generation (RAG) framework tailored for item cold-start recommendation, called ColdRAG. The key idea is to equip a large language model (LLM) with external knowledge, represented as a graph, to guide multi-hop reasoning beyond its internal knowledge from pre-trained data. This combination enables both semantic reasoning and evidence-grounded generation for recommending cold-start items. Rather than relying solely on internal LLM knowledge or static prompts, our approach dynamically builds a knowledge graph from item metadata and engages the LLM in structured reasoning and response composition based on retrieved information. Figure 1 illustrates the pipeline with four major stages: (1) Item Profile Generation, (2) Knowledge Graph Construction, (3) Multi-hop Reasoning-based Retrieval, and (4) Recommendation with Retrieval-augmented Generation.
3.1. Problem Statement
Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. Each user $u\in\mathcal{U}$ has an interaction history $H_{u}=[i_{1},i_{2},\dots,i_{n-1}]$, and the goal is to recommend the next item $i_{n}$ that the user is likely to interact with. In the cold-start setting, the target item $i_{n}$ is assumed to be previously unseen or rarely interacted with (i.e., a cold item). Our objective is to generate a top-$k$ recommendation list $\hat{\mathcal{Y}}_{u}\subseteq\mathcal{I}$ using LLM-guided reasoning over a dynamically constructed knowledge graph. This graph is built from item metadata and structured domain knowledge, enabling ColdRAG to retrieve and reason over relevant items without relying on historical interaction frequency or pre-filtered candidates.
3.2. Item Profile Generation
In real-world cold-start scenarios, item metadata is often sparse, incomplete, or lacks descriptive richness, making it difficult to extract sufficient semantic signals for recommendation. In addition, available metadata often contains noisy, redundant, or irrelevant content that may hinder effective reasoning. To address these challenges, we use a LLM to both enrich the metadata with knowledge-grounded content and distill it into a focused, natural language summary. The resulting item profile integrates structured fields with relevant background knowledge inferred by the model, while filtering out extraneous information and preserving elements most indicative of user preferences.
Formally, for each item $i\in\mathcal{I}$, we define its structured metadata as:
$\mathbf{m}_{i}=(\text{title}_{i},\text{description}_{i},\text{attributes}_{i},\text{reviews}_{i})$
For example, the item titled Tomb Raider may include verbose and loosely structured metadata such as: a product description like “The first of the now-famous Tomb Raider games is a must-have for every gaming fan…,” attribute statements such as “Over 5,000 frames of animation on the main character, Lara, alone.” and “The first real 3-D interactive exploratory adventure,” and user reviews including comments like “This game has GREAT historical value.”.
Our approach leverages an LLM to construct a clean and coherent summary—item profile—that retains key content while filtering out irrelevant phrasing. To enable this, we formulate a structured prompt using the available metadata $\mathbf{m}_{i}$ (including title, description, attributes, and user reviews) and instruct the LLM to generate a concise, engaging, and informative paragraph with the prompt—for example: “You are an expert in gaming and consumer electronics. You are provided with item titles, descriptions, attributes, and user reviews. For each item, write a concise, engaging, and informative description.”. This design guides the model to focus on semantically rich elements aligned with user preferences, resulting in a fluent, context-aware profile that enhances downstream reasoning. The generation step is formalized as:
$P_{i}=\textit{LLM}(\textit{prompt}(\mathbf{m}_{i}))$
As shown in Figure 1, the generated item profile for Tomb Raider depicts the process to distill verbose and loosely structured metadata into a coherent summary. It highlights essential concepts such as action-adventure, Lara Croft, diverse landscapes, and cinematic storytelling—terms that are only implicitly present in the original metadata but are surfaced through knowledge-grounded rewriting. The result is a semantically rich and compact profile that captures the core themes relevant to user interest.
This item profile serves as the foundation for constructing the knowledge graph. Its structured narrative enables the accurate extraction of entities and relations, which form the nodes and edges of the graph. By abstracting noise and standardizing item-level semantics, the framework ensures that the resulting graph is well-suited for multi-hop reasoning—particularly in item cold-start scenario.
3.3. Knowledge Graph Construction
To enable structured reasoning and context-aware recommendation in the next stage, the framework transforms the enriched item profiles $P_{i}$ into a semantic graph representation. Specifically, it constructs a knowledge graph $\mathcal{G}=(\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ is the set of entities and $\mathcal{R}$ is the set of directed relationships between them. This transformation serves two main purposes. First, it organizes the unstructured content of the item profiles into a structured and entity-centric format, allowing the framework to extract clearly defined and semantically meaningful concepts. Second, it enables retrieval and reasoning over related information through multi-hop traversal, which supports more informed and controllable recommendation generation. Compared to conventional RAG approaches that operate on flat, paragraph-level text chunks, adopting a graph-based retrieval framework provides access to more precise concept-level relationships, ultimately leading to richer and more targeted retrieval outcomes.
To build the knowledge graph $\mathcal{G}$, we extract structured entities and their relationships from each item profile utilizing LLM. Each entity $e\in\mathcal{E}$ is represented as a tuple $(e_{\text{name}},e_{\text{type}},e_{\text{desc}},e_{\text{emb}})$. $e_{\text{name}}$ denotes the surface name of the entity—for example, ”Tomb Raider”. $e_{\text{type}}$ specifies the semantic category of the entity, such as item, genre, or feature; in this case, item. $e_{\text{desc}}$ provides a natural language explanation of the entity’s role or relevance—for instance, ”groundbreaking action-adventure game featuring the fearless …”. Finally, $e_{\text{emb}}\in\mathbb{R}^{d}$ is a dense vector embedding of the entity description ($e_{\text{desc}}$), computed using a pretrained language model (e.g., OpenAI’s text-embedding-3-small [^1] https://platform.openai.com/docs/models/text-embedding-3-small). These embeddings are stored in a vector index to support efficient similarity-based retrieval during the multi-hop reasoning stage.
Each relation $r\in\mathcal{R}$ is represented as a tuple $(r_{\text{name}},r_{\text{desc}},r_{\text{emb}})$, where $r_{\text{name}}=(e_{\text{src}},e_{\text{tgt}})$ identifies the pair of linked entities. For example, (”Tomb Raider”, ”Lara Croft”) indicates that the item is semantically connected to the character. The component $r_{\text{desc}}$ describes this relationship in natural language—for instance, ”Lara Croft is the main character driving the player’s experience in Tomb Raider. The iconic protagonist…”. The field $r_{\text{emb}}\in\mathbb{R}^{d}$ is a dense embedding of the relation description ($r_{\text{desc}}$), computed using the same embedding model. These relation embeddings support semantic filtering and scoring during graph traversal in the following module. Together, these structured and embedded representations form the foundation of the ColdRAG’s multi-hop reasoning capabilities. An example of this process is illustrated in Figure 1, where the item profile for Tomb Raider yields multiple typed entities. For each non-item entity, a relation is created toward the item node.
Representing the knowledge base as a graph offers several advantages in the context of RAG (Peng et al., 2024). Unlike flat document stores or unstructured corpora, a knowledge graph encodes structured, relational knowledge about each item and its associated concepts. This graph structure captures semantic relationships between entities—such as shared genres, features, or narrative roles—allowing the framework to model complex dependencies that are otherwise difficult to express. As illustrated in Figure 1, the item profile for Tomb Raider yields multiple typed entities, such as Lara Croft and action-adventure, each connected back to the item node through semantically meaningful relationships. By organizing information in this way, the framework can retrieve enriched, contextually relevant content and reason over multiple interconnected entities. The graph’s relational nature also supports controllable multi-hop traversal, enabling the framework to selectively gather supporting information that is most relevant to a given user query.
To support retrieval at inference stage, the constructed knowledge graph is persistently stored in a modular knowledge base. The graph topology—consisting of nodes and edges—is saved in structured files that record each entity’s name, type, and linked connections. Alongside this, the framework stores the corresponding textual descriptions of all entities and relations, which are used to construct prompts during retrieval. These descriptions provide semantics that enhance the utility of the graph. To enable semantic search, each textual description is embedded using a pretrained transformer-based embedding model optimized for similarity tasks. These vector representations are stored in a vector database that enables approximate nearest neighbor (ANN) retrieval based on cosine similarity (e.g., FAISS [^2] https://github.com/facebookresearch/faiss). Such frameworks are designed to scale to high-dimensional embedding spaces and support efficient semantic search. This hybrid storage architecture enables the retrieval of relevant knowledge both structurally (via graph traversal) and semantically (via embedding similarity), ensuring fast, contextually appropriate, and high-quality retrieval for zero-shot recommendation.
3.4. Multi-hop Reasoning-based Retrieval
Once the knowledge graph $\mathcal{G}$ is constructed and stored, the framework performs LLM-guided reasoning to retrieve item entities that are semantically aligned with the user’s preferences. For a given user $u\in\mathcal{U}$ with an interaction history $H_{u}=[i_{1},i_{2},\dots,i_{n-1}]$, the process begins by extracting the titles of previously interacted items as keyword anchors. These are embedded using the identical pretrained language model used during Knowledge Graph Construction and matched against stored entity embeddings to retrieve the most semantically similar node for each keyword. Specifically, for a keyword embedding $\vec{q}$, the framework retrieves the entity $e^{*}\in\mathcal{E}$ that maximizes cosine similarity:
$e^{*}=\arg\max_{e\in\mathcal{E}}\cos(\vec{q},\vec{e})=\arg\max_{e\in\mathcal{E}}\frac{\vec{q}\cdot\vec{e}}{\|\vec{q}\|\cdot\|\vec{e}\|}$
Starting from the initial node set $\mathcal{E}_{0}$, the model performs iterative graph traversal via one-hop expansions. At each step $t$, it collects all outward edges from the current frontier $\mathcal{F}_{t}$ and scores each edge’s natural language description $r_{\text{desc}}$ using the LLM:
$s_{r}=\textit{LLMScore}(r_{\text{desc}},H_{u})$
Only edges with a score above a predefined threshold $\lambda\in[0,1]$ are retained. Their target nodes form the next traversal frontier:
$\mathcal{F}_{t+1}=\{e^{\prime}\mid(e,e^{\prime},r_{\text{desc}})\in\mathcal{R},\ s_{r}\geq\lambda\}$
If a target node $e^{\prime}$ is of type item, it is first added to a temporary candidate pool denoted as $\widetilde{\mathcal{C}}_{u}$, and its corresponding entity and relation descriptions are added to the temporary contextual text set $\widetilde{\mathcal{T}}_{u}$. In addition, any node in the initial frontier $\mathcal{E}_{0}$ that is also of type item is included in $\widetilde{\mathcal{C}}_{u}$ and $\widetilde{\mathcal{T}}_{u}$ prior to the traversal. This graph traversal continues until the number of collected item-type nodes reaches a predefined maximum pool size $\theta_{\text{pool}}$ (e.g., 500), i.e.,
$|\widetilde{\mathcal{C}}_{u}|\geq\theta_{\text{pool}}.$
After forming the full pool $\widetilde{\mathcal{C}}_{u}$, the model ranks all item candidates by their edge relevance scores. The top $\theta_{\text{top}}$ items (e.g., 100) are selected as the final candidate set $\mathcal{C}_{u}\subseteq\widetilde{\mathcal{C}}_{u}$. The corresponding entity and relation descriptions associated with these items are then gathered to form the final contextual text block $\mathcal{T}_{u}\subseteq\widetilde{\mathcal{T}}_{u}$. This two-stage approach—first expanding to a broader semantic neighborhood and then pruning to the most relevant candidates—ensures that the resulting candidate set and contextual information are both comprehensive and highly aligned with user preferences.
As illustrated in Figure 1, starting from the node Tomb Raider, the traversal may pass through semantic neighbors like action- adventure or franchise reboot, reaching item entities such as Uncharted 1 and Lord of the Rings 2. These become recommendation candidates. Meanwhile, edge descriptions such as “Uncharted 1 is the first game in the Uncharted series that follows …” are collected as contextual texts to support downstream reasoning.
Unlike conventional LLM-based recommenders that rely on a predefined candidate list provided in the input prompt, this approach dynamically discovers relevant candidate items through semantic reasoning over the knowledge graph. It removes the need to manually supply a candidate subset in advance, which is often constrained by input length limitations and can restrict the coverage and diversity of the recommendation results. By enabling open-ended, dataset-grounded retrieval, the framework is able to consider a broader and more contextually relevant set of items tailored to the user’s interests.
The outputs of this stage are: (1) a candidate item set $\mathcal{C}_{u}\subseteq\mathcal{I}$, and (2) a contextual text block $\mathcal{T}_{u}$, composed of the descriptions of the entities and relations encountered during traversal. These are passed to the next stage for final recommendation generation.
Figure 2. Illustration of ColdRAG’s adaptability to item cold-start scenario.
3.5. Recommendation with Retrieval-augmented Generation
In the final stage, the framework generates a recommendation list using the outputs of the Multi-hop Reasoning-based Retrieval module: (1) the candidate item set $\mathcal{C}_{u}$, and (2) the contextual text block $\mathcal{T}_{u}$. These are used to construct the LLM prompts for response generation. The contextual text $\mathcal{T}_{u}$, composed of natural language descriptions of high-scoring relations and their associated entities retrieved from the knowledge graph, is provided as the system prompt. Each included description corresponds to a relation that was deemed relevant to the user’s query and leads to an item-typed entity, thereby offering semantic clues as to why that item is contextually appropriate for recommendation. In parallel, the candidate item set is injected into the original user query, forming the user prompt that communicates preference signals and constrains the output space to relevant items discovered during graph traversal.
The LLM takes both prompts simultaneously and generates a ranked list of item titles likely to match the user’s interests. Specifically, the contextual text block $\mathcal{T}_{u}$ is provided as the system prompt, while the user prompt $Q_{u}$ is constructed by injecting the candidate item set $\mathcal{C}_{u}$ into the original user query. This setup ensures that the model considers both the user’s preferences and relevant candidate items discovered through graph-based retrieval. The generation is constrained implicitly by the candidate set included in the user prompt, keeping outputs within the dataset. The model’s natural language response is then parsed to extract the top-$k$ recommendations:
$\hat{\mathcal{Y}}_{u}=\textit{ParseTopK}(\textit{LLM}(\mathcal{T}_{u},Q_{u}),k)$
where $\textit{LLM}(\mathcal{T}_{u},Q_{u})$ denotes the generation conditioned on the system and user prompts, and ParseTopK is a rule-based parser that selects the top-ranked item titles from the generated text. This design ensures that the recommendations are grounded in both user preferences and structured, graph-based context. By explicitly aligning generation with retrieved knowledge, the model produces more relevant and dataset-constrained results—especially beneficial in item cold-start scenarios where historical interaction data is sparse or unavailable.
Furthermore, ColdRAG framework offers strong adaptability to item cold-start scenarios. When a new item is introduced with its metadata, the framework can immediately generate its item profile, extract semantic entities and relations, and integrate them into the knowledge graph using the same pipeline applied to existing items. As illustrated in Figure 2, this allows the new item to be seamlessly inserted into the graph structure through its extracted concepts and connections. Once integrated, the item becomes eligible for graph-based reasoning and retrieval, enabling it to participate in recommendations without requiring any historical interactions or additional training. This flexibility ensures that the system remains robust and responsive in dynamic environments where new content is frequently introduced.
Table 2. Summary of dataset statistics after filtering.
Dataset #Interactions #Items #Users
Games 45,106 2,027 2,096
Toys 332,055 12,342 20,390
Office 233,738 6,107 15,302
Table 3. Statistics of the constructed knowledge graphs.
Dataset #Nodes #Edges Degree Density(%)
Games 15,048 29,023 3.86 0.0256
Toys 58,096 132,229 4.55 0.0078
Office 42,769 75,053 3.51 0.0082
Table 4. Comparison of Recall@10 and NDCG@10 across three datasets: Games, Toys, and Office. Our proposed models are highlighted in gray. For each column, the best result is highlighted in bold, and the second-best is underlined. The final row shows the improvement of our best-performing model compared to the best baseline per dataset. All results are averaged over 5 runs.
Model Games Toys Office
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
full- train UniSRec 1.33 0.54 1.48 0.71 1.41 0.66
CLCRec 5.71 2.98 2.75 1.40 3.29 1.97
TDRO 7.51 4.21 2.56 1.29 3.65 2.08
zero- shot LLM 3.48 1.52 1.32 0.60 3.16 1.34
LLMRank (S) 4.62 1.94 1.36 0.58 3.80 1.63
LLMRank (R) 8.86 4.25 1.13 0.72 4.20 2.23
LLMRank (I) 6.20 3.78 1.20 0.70 4.00 1.90
TaxRec 3.75 1.78 0.60 0.34 2.20 0.96
KALM4Rec 8.50 4.14 4.26 2.13 3.43 1.71
ColdRAG (ours) 12.37 4.38 5.40 2.29 8.60 3.26
Improvement +39.62% +3.06% +26.76% +7.51% +104.76% +46.19%
4. Experiments
We conduct comprehensive experiments to evaluate the effectiveness of ColdRAG in item cold-start recommendation. To guide our analysis, we organize our evaluation around the following five research questions:
  • RQ1: Does ColdRAG effectively address the item cold-start recommendation problem?
  • RQ2: How effective is the Knowledge Graph Construction module to provide knowledge grounding?
  • RQ3: How effective is the Multi-hop Reasoning-based Retrieval mechanism in retrieving preference-aligned items?
  • RQ4: Does ColdRAG exhibit improved generation stability in its recommendations?
  • RQ5: Does ColdRAG reduce hallucination by avoiding out-of-domain recommendations?
4.1. Experimental Setup
4.1.1. Datasets
We conduct experiments on three domains from the Amazon Review dataset (Ni et al., 2019) to evaluate the performance of ColdRAG in item cold-start recommendation scenarios: Games, Toys, and Office. These datasets span diverse product categories and interaction patterns, allowing us to assess the effectiveness of our method across various types of recommendation tasks. To ensure sufficient user and item coverage, we apply core filtering to remove sparse interactions. Specifically, we perform 15-core filtering on the Games dataset, and 10-core filtering on the Toys and Office datasets. This process retains only users and items with at least 15 or 10 interactions, respectively, resulting in well-connected and reliable data suitable for sequential recommendation tasks.
To simulate the cold-start setting, we identify the bottom 10% least frequently interacted items in each dataset as cold items. We follow a sequential recommendation setting, treating each user’s interaction history as a sequence. Specifically, we adopt the widely used leave-one-out evaluation (Sun et al., 2019; Hou et al., 2022), where the last item in each user sequence is held out as the target for prediction. From each dataset, we sample 500 user sequences that end with a cold item. For each sequence, the first n1n{-}1 items are used as input to the model, and the final $n$th item is used as the test target.
For full-train baselines, we use the 500 cold-item user sequences as the test set and treat all remaining user sequences in the dataset as the training set. This design allows us to fairly compare zero-shot and fully supervised (full-train) methods under consistent evaluation protocols, while ensuring that cold items are properly isolated at inference phase. The statistics of the datasets after preprocessing are summarized in Table 2. The knowledge graph statistics constructed from the item profiles are reported in Table 3, with a detailed breakdown of the entity and edge type distributions provided in Figure 5.
4.1.2. Baselines
We compare ColdRAG against the following baselines from both full-train and zero-shot recommendation paradigms:
  • UniSRec (Hou et al., 2022): A universal sequence representation learning model that leverages contrastive learning and item text encoders to model user behavior. We fine-tune the pre-trained UniSRec model on domain-specific data. This baseline serves as a strong representation based recommender for comparison.
  • CLCRec (Wei et al., 2021): A contrastive learning framework specifically designed for cold-start recommendation. It augments user-item interactions with feature-based positive sampling and contrastive objectives to better generalize to cold items.
  • TDRO (Lin et al., 2024): The cold-start recommender based on temporally and distributionally robust optimization. It explicitly addresses data distribution shifts for cold items by optimizing worst-case performance across dynamic environments.
  • LLM: A plain large language model baseline that takes the user’s interaction history and a randomly sampled candidate item list as input prompt, without any knowledge retrieval or structure-aware context.
  • LLMRank (Hou et al., 2024): A zero-shot ranking method where LLMs score item relevance from a candidate set using natural language prompts. We evaluate three variants: LLMRank (S) (sequential prompting), LLMRank (R) (recency-focused prompting), and LLMRank (I) (in-context prompting with exemplars).
  • TaxRec (Liang et al., 2025): A taxonomy-guided LLM approach for zero-shot recommendation. It builds and uses a taxonomy to discover and rank relevant items. For fair comparison, we adopt a candidate list setting and apply taxonomy-based ranking over the candidates.
  • KALM4Rec (Kieu et al., 2024): A recent zero-shot recommendation method that retrieves relevant item knowledge based on user query keywords and incorporates the retrieved context into the prompt for cold-start generation.
The first three methods are full-train baselines, which involve training or fine-tuning model parameters on the target dataset. UniSRec is a pre-trained sequence representation model that is further fine-tuned for the downstream task. CLCRec and TDRO are well-established cold-start recommenders, with TDRO in particular achieving state-of-the-art performance in recent evaluations. The remaining four are zero-shot baselines that do not require model training or fine-tuning. LLM and LLMRank serve as representative LLM-based recommenders. TaxRec and KALM4Rec are recent zero-shot methods that address the cold-start problem using retrieval-augmented prompting and structured query understanding. These methods are selected to benchmark ColdRAG against a diverse set of architectures and modeling assumptions.
4.1.3. Evaluation Metrics
To evaluate the performance of the recommendation, we use two metrics widely adopted in cold-start recommendation tasks: Recall@$k$ and NDCG@$k$. Following prior works (Hou et al., 2022; Wei et al., 2021; Liang et al., 2025), we report results at $k=10$.
4.1.4. Implementation Details
We use the Azure OpenAI API to access all LLM components used in ColdRAG and other zero-shot baselines. The backbone LLM is gpt-4o-mini [^3] https://platform.openai.com/docs/models/gpt-4o-mini, and for semantic similarity, we use OpenAI’s text-embedding-3-small model to encode entity and relation descriptions. Vector embeddings are indexed using FAISS to support fast approximate nearest neighbor retrieval. In the Multi-hop Reasoning-based Retrieval module, we set the edge scoring threshold to 7 and the minimum candidate item set size to 500. For the zero-shot baselines other than ColdRAG (e.g., LLM, LLMRank, and TaxRec), the candidate list is composed of 100 items randomly sampled from the dataset. We conduct all experiments with 5 independent runs and report the average performance across runs to ensure reliability and stability of the results.
4.2. Results
4.2.1. Overall Performance (RQ1)
Table 4 presents the performance comparison across all baseline models and our proposed method on three datasets: Games, Toys, and Office. As shown, ColdRAG consistently achieves the highest performance across all datasets and metrics, demonstrating its effectiveness in addressing the item cold-start challenge. The model outperforms both full-train and zero-shot baselines, achieving an average improvement of +39.62% Recall@10 on Games, +26.76% on Toys, and +104.76% on Office compared to the best-performing baseline in each domain. Notably, the improvements in Recall are consistently more pronounced than those in NDCG. This suggests that ColdRAG’s primary strength lies in its multi-hop reasoning-based retrieval, which excels at identifying and including the correct item within the top candidates.
For the full-train baselines, TDRO achieves overall high performance by leveraging temporally and distributionally robust optimization to address data distribution shifts. However, it is still outperformed by ColdRAG across all domains. For instance, in the Office domain, ColdRAG achieves a Recall@10 score that is substantially higher than TDRO, indicating its superior ability to recommend cold-start items under sparse conditions. These results suggest that even advanced fully supervised models face limitations in capturing item-level semantics in cold-start setting, whereas ColdRAG’s retrieval-augmented, knowledge-grounded design enables more effective and adaptive recommendation.
Among the zero-shot baselines, the plain LLM and LLMRank variants (LLMRank (S), (R), and (I)) demonstrate varying degrees of effectiveness depending on how they prompt the LLM, yet all rely on prompts and pre-defined candidate lists. This limitation hinders their ability to adaptively explore semantic relationships beyond what is explicitly given in the prompt. KALM4Rec performs better by leveraging keyword-based retrieval to enrich prompts with relevant context. However, ColdRAG surpasses all these approaches by integrating structured, knowledge-grounded reasoning through multi-hop traversal over a dynamically constructed knowledge graph. This allows it to retrieve richer and more semantically aligned item context, leading to superior recommendation quality.
In summary, the strong and consistent performance of ColdRAG across both full-train and zero-shot baselines highlights the effectiveness of integrating structured external knowledge and graph-based reasoning into LLM-based recommendation. These results affirm the strength of ColdRAG’s design and its promise as a robust solution for zero-shot recommendation in item cold-start scenario.
Figure 3. Performance comparison of ColdRAG variants across three domains. “w/o G,M” refers to the plain LLM without knowledge grounding and reasoning. “w/o M” includes knowledge grounding but removes multi-hop reasoning.
4.2.2. Ablation Study (RQ2 & RQ3)
To assess the contribution of ColdRAG’s main modules, we perform an ablation study, evaluating three variants: (1) w/o G,M, a plain LLM baseline without knowledge grounding (G) or multi-hop reasoning (M)—relying solely on user history and randomly sampled candidates; (2) w/o M, a ColdRAG variant that includes the knowledge graph (G) for entity retrieval but removes the reasoning component, retrieving only items matched directly to the user history without multi-hop traversal (M); and (3) the full ColdRAG, which leverages both G and M to construct prompts through dynamic graph traversal.
As shown in the Figure 3, performance improves steadily from w/o G,M to ColdRAG, particularly in the Games and Toys domains. This trend indicates that knowledge graph construction (G) introduces semantically enriched signals that support more informed generation, while multi-hop reasoning (M) enables flexible exploration beyond local item matches—resulting in more contextually relevant recommendations. In the Office domain, we observe that the w/o M variant slightly underperforms the w/o G,M baseline. This suggests that in domains with sparse or noisy metadata, naively incorporating external knowledge without proper reasoning may introduce irrelevant or distracting context. However, once reasoning is enabled in the full ColdRAG, performance rebounds significantly. This demonstrates that the multi-hop reasoning component plays a crucial role in refining and contextualizing knowledge retrieval, selectively filtering useful information while avoiding noise.
In summary, ColdRAG’s Knowledge Graph Construction and Multi-hop Reasoning-based Retrieval modules are both effective and complementary. The former enriches semantic context, while the latter selectively retrieves relevant signals. Together, they enable effective recommendations in cold-start scenarios.
Figure 4. (a) Box plot comparison of Recall@10 across five runs for zero-shot baselines on the Games dataset. (b) Out-of-domain generation rate comparison on the Games dataset.
4.2.3. Analysis on Generation Stability (RQ4)
LLMs are known to exhibit generation inconsistency—producing varied outputs for the same input depending on minor prompt changes or decoding randomness (Elazar et al., 2021). This inconsistency can undermine reliability in recommendation settings, where stable outputs are crucial for user trust. To investigate ColdRAG’s robustness, we assess generation stability by measuring the variance in recommendation quality across five runs. Figure 4 (a) displays the performance distributions of Recall@10 on the Games dataset using box plots for each zero-shot model, offering insight into output consistency.
ColdRAG consistently exhibits both high performance and low variance across runs, highlighting its ability to produce stable recommendation outputs. In contrast, other zero-shot baselines show wider distributions, indicating greater sensitivity to randomness during generation. Notably, we exclude KALM4Rec for fair comparison, as it prompts the LLM to generate item identifiers instead of natural language-formed item titles, making its output format incompatible with our evaluation setup. This stability can be attributed to ColdRAG’s use of structured, knowledge-grounded retrieval via multi-hop reasoning over a semantic graph. Rather than relying solely on raw prompt inputs, ColdRAG constructs contextual signals that are consistent and semantically relevant, enabling the LLM to make more reliable decisions. These findings suggest that ColdRAG not only outperforms other models in accuracy but also provides a more deterministic and controllable recommendation process, which is particularly important in production environments where reproducibility matters.
4.2.4. Analysis on Hallucination (RQ5)
A persistent issue in LLM-based recommendation is hallucination—the generation of invalid item titles that are absent from the target dataset. Such out-of-domain outputs reduce the reliability and usability of the recommendations in real-world applications. To evaluate ColdRAG’s effectiveness in mitigating hallucination, we analyze the proportion of out-of-domain items generated on the Games dataset, as shown in Figure 4. For a fair comparison, all baselines—LLM, LLMRank(S), LLMRank(R), LLMRank(I), and TaxRec—are provided with explicitly constructed candidate lists in prompts for controlled generation.
Despite this constraint, these models still produce a substantial number of out-of-domain items, with hallucination rates ranging from 5.82% to 9.89%. In contrast, ColdRAG achieves a significantly lower hallucination rate of just 3.15%. This indicates that even with access to candidate lists, prompt-based control alone is insufficient to prevent hallucination. ColdRAG eliminates the need for such predefined candidate lists by leveraging a dynamically constructed knowledge graph for retrieval. Its LLM-guided multi-hop reasoning mechanism allows it to retrieve semantically relevant, in-domain content grounded in structured knowledge. This grounding substantially reduces hallucination and improves the controllability and trustworthiness of the generated recommendations.
Figure 5. (a) Pie chart of the distribution of entity (node) types in the KG. (b) Heatmap of the number of relations (edges) between the entity types.
4.2.5. Structure of Knowledge Graph
To provide better understanding about the structure of the generated KG constructed in Knowledge Graph Construction, we analyzed the composition of its nodes and edges in the Games domain. Figure 5 (a) illustrates the distribution of node types. The KG is predominantly composed of item nodes, which represent half of all entities (50%), and feature nodes, making up another significant portion (26%). Other entity types such as target user (9%), etc (6%), setting (5%), and genre (4%) constitute the remainder, providing rich, multi-faceted context for the items. This distribution highlights that the KG is naturally centered around the items themselves, as all other non-item entities are components generated to describe and explain the items. The feature category is the second most abundant because it serves as a broad, versatile concept that can capture a wider array of attributes for an item.
The heatmap in Figure 5 (b) details the connectivity between these different node types, showing the number of edges. A key observation is the dense connection between item and feature nodes, with 15,102 edges, confirming that items are primarily defined by a rich set of features within the graph. Similarly, the substantial number of edges between item and target user nodes (4,928) indicates that the graph effectively captures relationships relevant to user preferences. The connections between items and other concepts like genre (2,542) and setting (2,735) further enrich the semantic network, enabling more nuanced, multi-hop reasoning during the recommendation process. This interconnected structure demonstrates that the KG successfully models the complex relationships between items, their attributes, and user contexts, forming a solid foundation for the retrieval and reasoning stages of ColdRAG.
5. Conclusion
In this work, we introduced ColdRAG, a retrieval-augmented generation framework that addresses the item cold-start problem by integrating LLM-based reasoning with a dynamically constructed knowledge graph. Unlike typical graph-RAG systems that rely on simple text similarity, ColdRAG employs LLM-guided, multi-hop reasoning to retrieve evidence tailored to user preferences. This targeted approach allows the framework to handle sparse metadata, enhance recommendation quality, and reduce out-of-domain generation without requiring task-specific training. Extensive experiments validate that ColdRAG consistently outperforms state-of-the-art baselines in accuracy, stability, and reliability, offering a practical zero-shot solution for cold-start recommendation.